# Measurement of the mean excitation energy of liquid argon

### v1

@@ 5.4.5 Other material adjustments
The vessel was filled with liquid argon with a total depth of 16 cm. Data was taken between two and four hours after the fill. Previous tests without beam showed that the loss rate of liquid argon is 0.3 cm per hour. Each beam pulse adds additional heat, but this effect increases the evaporation rate by only 0.1%. Therefore, we estimate that the argon level varied from 15.4 cm to 14.8 cm during the experimental run. Monte Carlo simulations used for the fit assumed a depth of 12 cm, but we find no significant differences in simulation results for depths between 11 cm and 16 cm. Even at 11 cm, there is continuous argon for all straight-line paths from the beam pipe to the scintillator, but small modification in transmission begin to occur because protons which are scattered upwards into the air have no significant chance of being scattered back down towards the scintillator.
@@ 5.7.4 Straggling
To test if Geant4’s model for range straggling — the variation in range for particles with same initial energy — biases the result, the experiment was first simulated with the extreme case of no straggling, such that the only significant variations in protons’ energy reaching the scintillator were from differing path lengths due to multiple scattering and from the energy spread in the initial beam. Using these simulations to analyze the data instead of the standard simulations resulted in an downwards shift of 4.7 eV in the inferred I-value. As a less extreme case, the Geant4 UniversalFluctuations model was substituted for more accurate UrbanFluctuations model (which is the default for both Opt0 and Opt4/EMZ), which resulted in a downwards shift of 0.5 eV. From this, we conclude that further improvements in the straggling model are unlikely to result in shifts of more than half this amount and adopt 0.3 eV as the systematic error from the straggling model, which is negligible compared to the uncertainty from multiple scattering.
@@ 5.8.1 Vessel alignment
The vessel’s vertical position was determined to 0.5 cm. Similarly to the case of the liquid argon depth, above, there is some slight sensitivity to the vertical position even if all straight-line proton paths lead entirely through argon because protons can scatter from the argon into the aluminum vessel and then back into the argon. Simulation showed that the effect on the I-value is only ±$\pm 0.2$ eV.
@@ 6 Conclusions
Figure 8: Combination of prior information with these results. The gray band shows the result combining all past and present information. The present result for the on-axis and off-axis scintillators are shown on the right, with statistical and total errors.
The present results are shown in figure 8 in comparison to the previous evaluations in Ref. [4] for gaseous and liquid argon. The overall result from this measurement combines the on-axis and off-axis scintillator results, assuming full correlation of systematic uncertainties, to arrive at $(205\pm 4)$ eV.
To make our overall recommendation for the I-value of liquid argon, we combine our previous evaluation for liquid argon [4] with the present results. No significant sources of correlated uncertainties were identified between the present measurement and the previous evaluation from Ref. [4], which relied primarily on oscillator strength distributions (i.e. photoabsorption cross sections) and an estimate of the phase effect using previous measurements several other substances in gaseous and condensed forms. Secondary contributions to the previous evaluations come from consideration of periodic trends, a Hartree-Fock wave function calculation, and a stopping power measurement using natural alpha particles in gaseous argon. Even smaller contributions are from four other stopping power measurements. Some correlation could exist in the treatment of multiple Coulomb scattering in the stopping power measurements. However, this group of measurements collectively make so little contribution that the result is unchanged to the last reported decimal place if they are discounted entirely. Therefore we treat the evaluation and this measurement as independent. To avoid loss of accuracy, we carried all digits through the calculation, using $(196.8\pm 6.5)$ eV for the previous evaluation and $(204.9\pm 3.6)$ eV for the present meeasurement. The combined recommendation is $(203.0\pm 3.2)$ eV.
Using our recommended value rather than the old values of 188 eV or 187 eV shifts energy reconstruction in liquid argon TPCs used for GeV-scale neutrinos by around 0.6%, with the exact shift depending on the analysis method. The inferred value of $\Delta m^{2}_{32}$ is directly proportional to reconstructed energy. As DUNE projects a sensitivity to this parameter of around 0.5% (depending on exposure) [27], is it important that a reliable value and uncertainty for the mean excitation energy of liquid argon be used.

### v2

@@ 5.3.5 Other material adjustments
The vessel was filled with liquid argon with a total depth of 16 cm. Data was taken between two and four hours after the fill. Previous tests without beam showed that the loss rate of liquid argon is 0.3 cm per hour. Each beam pulse adds additional heat, but this effect increases the evaporation rate by only 0.1%. Therefore, we estimate that the argon level varied from 15.4 cm to 14.8 cm during the experimental run. Monte Carlo simulations used for the fit assumed a depth of 12 cm, but we find no significant differences in simulation results for depths between 11 cm and 16 cm. Even at 11 cm, there is continuous argon for all straight-line paths from the beam pipe to the scintillator, but small modifications in transmission begin to occur because protons which are scattered upwards into the air have no significant chance of being scattered back down towards the scintillator.
@@ 5.4 Multiple Coulomb scattering
5.5 Monte Carlo code consistency
[…]
[width=0.5]fig-mars-geant.pdf
The other Monte Carlo codes predict slightly shorter or longer proton ranges than Geant4 given the same I-value, meaning that analyzing the data using them results in, respectively, higher or lower I-values. FLUKA and mars each give a result 1.6 eV higher than Geant4, while the answer from PHITS was 1.0 eV lower. As is visible from the plot, the largest difference between the codes is on the left side, before the peak, although it is the falling edge on the right side which gives the range. The differences between codes to the left of the peak are much smaller than the detector effects discussed in section 5.1.
@@ 5.7.4 Straggling
To test if Geant4’s model for range straggling — the variation in range for particles with same initial energy — biases the result, the experiment was first simulated with the extreme case of no straggling, such that the only significant variations in protons’ energy reaching the scintillator were from differing path lengths due to multiple scattering and from the energy spread in the initial beam. Using these simulations to analyze the data instead of the standard simulations resulted in an downwards shift of 4.7 eV in the inferred I-value. As a less extreme case, the Geant4 UniversalFluctuations model was substituted for more accurate UrbanFluctuations model (which is the default for both Opt0 and Opt4/EMZ), which resulted in a downwards shift of 0.5 eV. From this, we conclude that further improvements in the straggling model are unlikely to result in shifts of more than half this amount and adopt 0.3 eV as the systematic error from the straggling model.
5.7.5 Further numerical approximations
@@ 5.8.1 Vessel alignment
The vessel’s vertical position was determined to 0.5 cm. Similarly to the case of the liquid argon depth, above, there is some slight sensitivity to the vertical position even if all straight-line proton paths lead entirely through argon because protons can scatter from the argon into the aluminum vessel and then back into the argon. As the argon-aluminum boundary is much closer to the beam center than the argon-air boundary, there is more possibility of this effect being significant. Simulation showed that the effect on the I-value is only ±$\pm 0.2$ eV.
@@ 6 Conclusions
[…]
[width=0.4816]fig-evidence.pdf
Figure 9: Combination of prior information with these results. The shaded band shows the result combining all past and present information. The present result for the on-axis and off-axis scintillators are shown on the right, with statistical and total errors.
The present results are shown in figure 9 in comparison to the previous evaluations in Ref. [4] for gaseous and liquid argon. The overall result from this measurement combines the on-axis and off-axis scintillator results, assuming full correlation of systematic uncertainties, to arrive at $199^{+4}_{-5}$ eV.
To make our overall recommendation for the I-value of liquid argon, we combine our previous evaluation for liquid argon [4] with the present results. No significant sources of correlated uncertainties were identified between the present measurement and the previous evaluation from Ref. [4], which relied primarily on oscillator strength distributions (i.e. photoabsorption cross sections) and an estimate of the phase effect using previous measurements several other substances in gaseous and condensed forms. Secondary contributions to the previous evaluations come from consideration of periodic trends, a Hartree-Fock wave function calculation, and a stopping power measurement using natural alpha particles in gaseous argon. Even smaller contributions are from four other stopping power measurements. Some correlation could exist in the treatment of multiple Coulomb scattering in the stopping power measurements. However, this group of measurements collectively make so little contribution that the result is unchanged to the last reported decimal place if they are discounted entirely. Therefore we treat the evaluation and this measurement as independent. To avoid loss of accuracy, we carried all digits through the calculation, using $(196.8\pm 6.5)$ eV for the previous evaluation and $199.2^{+4.1}_{-5.0}$ eV for the present measurement. The combined recommendation is $(198\pm 4)$ eV.
All uncertainties have heretofore been given to $1\sigma$ (68% CL). At higher CL, the bounds tighten more quickly than the Gaussian case. At $2\sigma$ (95% CL), our result is $(199\pm 8)$ eV and our overall evaluation is $(198\pm 7)$ eV. Finally, at $3\sigma$ (99.7% CL), our result is $199^{+11}_{-12}$ eV and our evaluation is $(198\pm 10)$ eV. We therefore exclude 188 eV at the $3\sigma$ level.
Using our recommended value rather than the old values of 188 eV or 187 eV shifts energy reconstruction in liquid argon TPCs used for GeV-scale neutrinos by around 0.4%, with the exact shift depending on the analysis method. For instance, the dE/dx of a minimum-ionizing particle is reduced from 1.508 MeV g-1 cm-2 to $(1.501\pm 0.002)$ MeV g-1 cm-2. The inferred value of $\Delta m^{2}_{32}$ is directly proportional to reconstructed energy. As DUNE projects a sensitivity to this parameter of around 0.5% (depending on exposure) [32], is it important that a reliable value and uncertainty for the mean excitation energy of liquid argon be used.